Solve and check your solution. Show all work. [(5x+3)/ 2] + (3/4) = (5/2)
Answer by tara0066(31)

[(5x+3)/2]+(3/4)=5/2 First, subtract your 3/4 to both sides
[(5x+3)/2]+(3/4)-(3/4)=5/2-3/4 This cancels out your +3/4
[(5x+3)/2]=7/4 Next, multiply both sides by 2
2[(5x+3)/2]=7/4(2) This cancels out your divided by two
5x+3=7/2 Then, subtract your 3 over
5x+3-3=7/2-3 This cancels your +3
5x=1/2 Finally, you can divide by 5
5x(1/5)=1/2(1/5) Or you can multiply by the reciprical
of 5 which is 1/5
x=1/10